nico is saving money for his college education he invests some money at 7% and $2200 less than that at 6% the invests produced a total of $245 interest in 1 yr how much did he invest at each rate
Answer by richwmiller(17219)

We don't know the total invested. We know the difference in the two accounts.
We know that the interest for the two accounts is $245
0.07*x+0.06*y=245
We know that the account at 7% has $2200 more.
x=2200+y
We substitute for x
0.07*(2200+y)+0.06*y=245
We multiply out
154+0.07y+0.06*y=245
We combine like terms.
0.13*y=91
Isolate y
y=$700 at 6%
x=2200+y
Calculate x
x=$2900 at 7%
Total invested $2900+$700=$3600
We check
0.07*2900+0.06*700=245
203+42=245
245=245
Since this statement is TRUE then it is ok